1. A generic solid, X, has a molar mass of 79.4 g/mol. In a constant-pressure calorimeter, 21.9 g of X is dissolved in 303 g of water at 23.00 °C. The temperature of the resulting solution rises to 25.00 °C. Assume the solution has the same specific heat as water, 4.184 J/(g·°C), and that there\'s negligible heat loss to the surroundings. How much heat was absorbed by the solution? How much heat was absorbed by the solution? What is the enthalpy of the reaction?
Answer #1

Formula

1. .....eq(1)

Q = heat absorbed by solution.

c = specific heat

m = mass of solution.

2.

Enthalpy of reaction    ....(2)

Molar enthalpy of reaction = ....(3)

where n = number of moles of reactant.

Given:-

c = 4.184 J/g °c

Mass of X = 21.9 g Mass of water = 303 g

Molar mass of X = 79.4 g/mol

m = (mass of X + mass of water) = (21.9 + 303) = 324.9 g

= (25 °C - 23 °C) = 2 °C

From eq(1) we have

2.

Enthalpy of reaction =

Molar enthlpy of reaction =